# A Quantum Annealing Protocol to Solve the Nuclear Shell Model

### v2

@@ Abstract
The nuclear shell model accurately describes the structure and dynamics of atomic nuclei. However, the exponential scaling of the basis size with the number of degrees of freedom hampers a direct numerical solution for heavy nuclei. In this work, we present a quantum annealing protocol to obtain nuclear ground states. We propose a tailored driver Hamiltonian that preserves a large gap and validate our approach in a dozen nuclei with basis sizes up to $10^{5}$ using classical simulations of the annealing evolution. We explore the relation between the spectral gap and the total time of the annealing protocol, assessing its accuracy by comparing the fidelity and energy relative error to classical benchmarks. While the nuclear Hamiltonian is non-local and thus challenging to implement in current setups, the estimated computational cost of our annealing protocol on quantum circuits is polynomial in the many-body basis size, paving the way to study heavier nuclei.
@@ 1 Introduction
Quantum computing, algorithms, and devices are promising tools for studying quantum many-body systems. Two distinct types of approaches have been used to simulate ground states. On one hand, variational quantum eigensolvers (VQEs) [29, 30, 31, 32, 33] should provide an advantage in the Noisy Intermediate-Scale Quantum era. On the other hand, quantum annealers [34, 35, 36] offer a more physics-inspired approach to quantum simulation. In either approach, simulations have been applied not only to condensed matter systems [37, 38, 39, 40], but also to hadron physics [41] and quantum chemistry [42, 43, 44, 45, 46].
@@
Quantum annealing (QA) is a method used to find the ground state of Hamiltonians through time evolution. Recently, QAs [63, 64] have been widely used to solve quadratic unconstrained binary optimization (QUBO) problems [65, 66]. With QA, one aims to find the solution of this classical NP-hard problem by encoding it in the ground state of an Ising Hamiltonian. This is achieved via time evolution; however, the small spectral gap of the time-dependent Hamiltonian hampers the method’s effectiveness. Several techniques have been developed to improve the robustness of the annealing, such as tailored driver Hamiltonians [67, 68] or adaptive time schedules [69]. Moreover, the application range for QA has expanded beyond QUBO problems to include the ground state of condensed matter systems [39, 40] and molecules [42, 70, 71, 46], yielding promising results.
@@ 4 Driver Hamiltonian
Figure 1: Left panel: matrix elements of the driver Hamiltonian $\hat{H}_{D}$ in the single-particle basis (nucleon modes) for 8Be in the $p$ shell. The bottom left panel indicates the ordering of the modes, where left (right) single-particle levels correspond to neutrons (protons). Right panel: after the annealing protocol is finished, the Hamiltonian $\hat{H}(t)$ becomes the target Hamiltonian $\hat{H}(\tau)=\hat{H}_{T}$, which is significantly non-local in the many-body basis, described in the bottom right panel.
@@
which confines the entire adiabatic evolution into the $M=0$ sector of Fock space. Finally, it is noteworthy that our choice for the initial state corresponds to the minimum energy Slater determinant in $\mathcal{B}_{0}$
@@
in all studied nuclei except 20O, 22Ne, 24Mg and 26Mg. A lower energy may be advantageous for the QA protocol, although it does not guarantee faster or more efficient convergence than other initial states [starasco_TFM, 84].
@@ 6 Computational Cost of the Quantum Annealing Protocol on a Digital Quantum Computer
in terms of the local operators, $\hat{h}_{l}(t_{k})$, using the staircase algorithm to exponentiate the matrices [95]. To estimate the number of CNOT gates as a function of the number of nucleon modes per species $D$, we note that the number of two-body terms in the Hamiltonian is $\propto D^{4}$. In the Jordan-Wigner mapping, each two-body term has an average Pauli string length of $\propto D/2$, and the number of CNOT gates required to exponentiate a Pauli operator of length $D/2$ is $D$ [95]. Therefore, the number of CNOTs for a unitary gate is
$N_{\rm CNOT}\propto D^{5}\,,$ (22)
@@
These numbers of CNOT gates—for a single timestep—are comparable to the total required for the convergence of the largest-dimension nucleus in the same valence space with ADAPT-VQE, $2\cdot 10^{3}$ and $\sim 10^{6}$, respectively [51]. However, the QA CNOTs grow polynomially with $D$. This is, if we assume a constant number of time steps for all $D$, our approach may scale better than ADAPT-VQE, which requires an exponentially growing number of CNOTs proportional to the number of many-body states [43, 51]. Hence, the polynomial scaling of the QA cost could offer an advantage for NSM calculations in heavier nuclei compared to the ADAPT-VQE method.
@@ 7 Conclusion
In this work, we present a QA protocol to obtain the ground states of atomic nuclei, suitable for quantum devices with high connectivity. In particular, we discuss a way to implement a driver Hamiltonian different from the ones used in the standard QA approach on spin systems. Our protocol is stable for all studied nuclei, both in the $p$ and $sd$ shells, encompassing systems with many-body basis dimensions up to $10^{5}$ states. Our QA approach achieves relative energy errors smaller than $O(10^{-4})$ in most cases, with fidelities greater than $0.99$. The best performance is consistently found when using longer timescales, which correspond to slower annealing evolutions. Nonetheless, we notice that the scaling of the accuracy with the timescale varies between nuclei, with better scaling for those where the driver Hamiltonian fully occupies nucleon subshells. Additionally, we observe that the final time required to reach a high-quality fidelity, $\tau^{*}$, depends on the level gap, $\Delta$, in agreement with the adiabatic condition, $\tau\propto\Delta^{-2}$.
@@
Due to the nonlocality of the NSM Hamiltonian, a direct implementation of the QA protocol using the Jordan-Wigner mapping on current QAs with low connectivity is not feasible. To compare this approach with state-of-the-art VQEs, we estimate the computational cost of the QA protocol on a digital quantum device. While our method is more computationally expensive than the ADAPT-VQE algorithm for nuclei in the $p$ and $sd$ shells, it scales polynomially with respect to the number of many-body basis states. For instance, for the nucleus with largest basis dimension we have studied, 28Si, the number of CNOT gates needed in our QA implementation is about one order of magnitude larger than with ADAPT-VQE. This suggests that QA could already offer advantages for heavier nuclei in the $pf$ shell.
This work establishes a foundation for using QA to study atomic nuclei. Future investigations could explore qubit mappings that reduce the complexity of Pauli strings or assess QA implementations while considering the limitations of real quantum annealers. Additionally, applying the method to heavier nuclei in the $pf$ shell warrants further study to optimize the annealing protocol.

### v3

@@ Abstract
The nuclear shell model accurately describes the structure and dynamics of atomic nuclei. However, the exponential scaling of the basis size with the number of degrees of freedom hampers a direct numerical solution for heavy nuclei. In this work, we present a quantum annealing protocol to obtain nuclear ground states. We propose a tailored driver Hamiltonian that preserves a large gap and validate our approach in a dozen nuclei with basis sizes up to $10^{5}$ using classical simulations of the annealing evolution. We explore the relation between the spectral gap and the total time of the annealing protocol, assessing its accuracy by comparing the fidelity and energy relative error to classical benchmarks. While the nuclear Hamiltonian is non-local and thus challenging to implement in current setups, the estimated computational cost of our annealing protocol on quantum circuits is polynomial in the single-particle basis size, paving the way to study heavier nuclei.
@@ 1 Introduction
Quantum computing, algorithms, and devices are promising tools for studying quantum many-body systems. Two distinct types of approaches have been used to simulate ground states. On one hand, variational quantum eigensolvers (VQEs) [29, 30, 31, 32, 33] should provide an advantage in the Noisy Intermediate-Scale Quantum era. On the other hand, quantum annealers [34, 35, 36] offer a more physics-inspired approach to quantum simulation. In either approach, simulations have been applied not only to condensed matter systems [37, 38, 39, 40], but also to hadron and nuclear physics [41, 42] and quantum chemistry [43, 44, 45, 46, 47].
@@
Quantum annealing (QA) is a method used to find the ground state of Hamiltonians through time evolution. Recently, QAs [64, 65] have been widely used to solve quadratic unconstrained binary optimization (QUBO) problems [66, 67]. With QA, one aims to find the solution of this classical NP-hard problem by encoding it in the ground state of an Ising Hamiltonian. This is achieved via time evolution; however, the small spectral gap of the time-dependent Hamiltonian hampers the method’s effectiveness. Several techniques have been developed to improve the robustness of the annealing, such as tailored driver Hamiltonians [68, 69] or adaptive time schedules [70]. Several studies have been also focused on both the criteria of convergence [71, 72, 73] and the speed-up of the QA protocol via counteradiabatic drivings [74, 75]. Moreover, the application range for QA has expanded beyond QUBO problems to include the ground state of condensed matter systems [39, 40] and molecules [43, 76, 77, 47], yielding promising results.
@@ 4 Driver Hamiltonian
Figure 1: Left panel: matrix elements of the driver Hamiltonian $\hat{H}_{D}$ in the single-particle basis (nucleon modes) for 8Be in the $p$ shell. The bottom panel indicates the ordering of the modes, where left (right) single-particle levels correspond to neutrons (protons). Right panel: after the annealing protocol is finished, the Hamiltonian $\hat{H}(t)$ becomes the target Hamiltonian $\hat{H}(\tau)=\hat{H}_{T}$, which is significantly non-local in the many-body basis.
@@
which confines the entire adiabatic evolution into the $M=0$ sector of Fock space. Finally, it is noteworthy that our choice for the initial state corresponds to the minimum energy Slater determinant in $\mathcal{B}_{0}=\{\ket{s}\in\mathcal{B}\,\,|\,\,M\ket{s}=0\}$, where $\mathcal{B}$ indicates the many-body basis,
@@
in all studied nuclei except 20O, 22Ne, 24Mg and 26Mg. A lower energy closer to the ground-state energy of the target Hamiltonian may be advantageous for a potential QA realization, although we emphasize that it does not guarantee faster or more efficient convergence than other initial states [90, 91].
@@ 5 Simulations of Nuclei with a Quantum Annealing Protocol
Finally, we expect these main features of the QA protocol to be robust with respect to an enlargement of the Hilbert space for the same nucleus, which should mostly affect higher-energy excited states leaving the energy gap relatively stable. Likewise, we also anticipate similar results from using different nuclear Hamiltonians, as long as they capture well the experimental low-energy structure of the nuclei studied in this work. Indeed, we already explore two different sets of nuclear interactions for $p$- and $sd$-shell nuclei. These are phenomenological in nature, but the systematics of our results across the different mass numbers suggest that our results do not depend significantly on the nuclear interaction used.
@@ 6 Computational Cost of the Quantum Annealing Protocol on a Digital Quantum Computer
in terms of the local operators, $\hat{h}_{l}(t_{k})$, using the staircase algorithm to exponentiate the matrices [102]. To estimate the number of CNOT gates as a function of the number of nucleon modes per species $D$, we note that the number of two-body terms in the Hamiltonian is $\propto D^{4}$. In the Jordan-Wigner mapping, each two-body term has an average Pauli string length of $\propto D/2$, and the number of CNOT gates required to exponentiate a Pauli operator of length $D/2$ is $(D-2)$ [102]. Therefore, the number of CNOTs for a unitary gate is
$N_{\rm CNOT}\propto D^{4}(D-2)\,,$ (22)
@@
These numbers of CNOT gates—for a single timestep—are comparable to the total required for the convergence of the largest-dimension nucleus in the same valence space with ADAPT-VQE, $2\cdot 10^{3}$ and $\sim 10^{6}$, respectively [52]. However, the QA CNOTs grow polynomially with $D$. In addition to that, we expect that the number of time steps also scales polynomially with $D$, since $\tau\propto\Delta^{-2}$, and $\Delta$ approaches the energy gap between the ground state and the first resonance state in $H_{T}$, that for nuclei scales polynomially with $D$ [13, 14]. In this perspective, our approach may scale better than ADAPT-VQE, which requires an exponentially growing number of CNOTs proportional to the number of many-body states [44, 52]. Hence, the polynomial scaling of the QA cost could offer an advantage for NSM calculations in heavier nuclei compared to the ADAPT-VQE method.
@@ 7 Conclusion
We present a QA protocol to obtain the ground states of atomic nuclei, suitable for quantum devices with high connectivity. Our work is, to our knowledge, the first application of quantum adiabatic computing to address a nuclear structure problem, enabling a clear synergy between two so-far disconnected research areas. In particular, we discuss a way to implement a driver Hamiltonian different from the ones used in the standard QA approach on spin systems. Our protocol is stable for all studied nuclei, both in the $p$ and $sd$ shells, encompassing systems with many-body basis dimensions up to $10^{5}$ states. Our QA approach achieves relative energy errors smaller than $O(10^{-4})$ in most cases, with fidelities greater than $0.99$. The best performance is consistently found when using longer timescales, which correspond to slower annealing evolutions. Nonetheless, we notice that the scaling of the accuracy with the timescale varies between nuclei, with better scaling for those where the driver Hamiltonian fully occupies nucleon subshells. Additionally, we observe that the final time required to reach a high-quality fidelity, $\tau^{*}$, depends on the level gap, $\Delta$, in agreement with the adiabatic condition, $\tau\propto\Delta^{-2}$.
@@
Due to the nonlocality of the NSM Hamiltonian, a direct implementation of the QA protocol using the Jordan-Wigner mapping on current QAs with low connectivity is not feasible. To compare this approach with state-of-the-art VQEs, we estimate the computational cost of the QA protocol on a digital quantum device. While our method is more computationally expensive than the ADAPT-VQE algorithm for nuclei in the $p$ and $sd$ shells, it scales polynomially with respect to the number of nucleon modes because the scaling of the energy gap of the target Hamiltonian is also polynomial. For instance, for the nucleus with largest basis dimension we have studied, 28Si, the number of CNOT gates needed in our QA implementation is about one order of magnitude larger than with ADAPT-VQE. This suggests that QA could already offer advantages for heavier nuclei in the $pf$ shell.
This work establishes a foundation for using QA to study atomic nuclei. The quantum adiabatic framework for nuclear structure presented here opens several new potential research avenues at the interface between these two research areas. Future investigations could explore criteria of convergence of the QA protocol without recurring to classical simulations, based on the form of the scheduler or the structure of the driver Hamiltonian [71, 72, 73]. Moreover, it is possible to explore ways to speed up the QA protocol by using counteradiabatic drivings [74, 75], studying the structure of the adiabatic gauge potential [108, 98] or using quantum optimal control techniques on the scheduler [70]. Future studies can also be focused on the qubit mappings that reduce the complexity of Pauli strings or assess QA implementations while considering the limitations of real quantum annealers. Additionally, applying the method to heavier nuclei in the $pf$ shell warrants further study to optimize the annealing protocol.